A manager wishes to build a 3-sigma range chart for a process. The sample size is five, the mean of sample means is 16.01, and the average range is 5.3. From Table S6.1, the appropriate value of D3 is 0, and D4 is 2.115. What are the UCL and LCL, respectively, for this range chart?

   

33.9 and 11.2

   

6.3 and 0

   

11.2 and 0

   

33.9 and 0

which is the correct answer?
Answer #1

Average Range R bar = 5.3

UCL = D4 Rbar = 2.115 * 5.3 = 11.2095

LCL = D3 Rbar = 0 * 5.3 = 0

Hence, the answer is Option C ( 11.2 and 0)